Two events and have probabilities 0.25 and 0.50, respectively. The probability that both and occur simultaneously is 0.14. Then, the probability that neither nor occurs, is
- A.0.39
- B.0.25
- C.0.11
- D.None of these
Show answer and solution
Correct option: A
Solution
P(A∪B) = P(A) + P(B) - P(A∩B) = 0.25 + 0.50 - 0.14 = 0.61.
P(neither A nor B) = 1 - P(A∪B) = 1 - 0.61 = 0.39.
P(neither A nor B) = 1 - P(A∪B) = 1 - 0.61 = 0.39.
[1980]